20ml each of O2 and H2 and 10ml of CO are exploded in an enclosure. What will be the volume and composition of the mixture of gases when they are cooled to room temperature?
1
Expert's answer
2014-02-03T07:28:24-0500
The volume of gas is directly proportional to number of moles. Hence we can write the equation for this process:
2xO2 + 2xH2 + xCO = x/2O2 + 2xH2O + xCO2
Where x is the number of moles in 10 ml of gas.
If the gas volumes in the initial mixture was given for the same conditions as the final, then we can calculate the final mixture composition:
V(O2) = x/2 = 5 ml;
V(H2O) = 2x = 20 ml (if the water is gasous. If not, its volume is negligible)
V(CO2) = x = 10 ml